OpenStudy (whovianchick):

PRE-CALCULUS HELP WILL FAN AND MEDAL :))))) Compare the graphs of the logarithmic functions f(x)=log_7 (x) and g(x)=log_4 (x). For what values of x is f=g, f>g, and f

OpenStudy (lexi):

f=g because you have to calculate the total of how ur going to start off.

satellite73 (satellite73):

hint \[\log_7(x)=\frac{\ln(x)}{\ln(7)}\]and \[\log_4(x)=\frac{\ln(x)}{\ln(4)}\]via "change of base"

OpenStudy (whovianchick):

So would you set them =, <, and > to each other and solve?

satellite73 (satellite73):

course you could always graph them with the computer you are on

OpenStudy (whovianchick):

True. How would I explain "how do you know" then lol

OpenStudy (tgstudios):

omg @satellite73 the 100 percenter... is that possible on openstudy?

satellite73 (satellite73):

what do you think is bigger, \[\log_4(16)\] or \[\log_7(16)\]?

OpenStudy (amorfide):

All of my years doing logarithms and I did not know that log_7(x) = lnx/ln7 thanks satellite

satellite73 (satellite73):

yw

OpenStudy (whovianchick):

Log_4 would be bigger

satellite73 (satellite73):

yes of course, now for what value of x do you think they are equal?

OpenStudy (whovianchick):

Well, 1 because that's where they intersect on the graph.

satellite73 (satellite73):

this is one number that gives the same answer no matter what the base of the log is do you know that number?

OpenStudy (whovianchick):

oh. No lol sorry

satellite73 (satellite73):

oh yes, what you said, 1

OpenStudy (whovianchick):

Oh ok

satellite73 (satellite73):

\[\log_b(1)=0\] which is another way of saying \[b^0=1\]

satellite73 (satellite73):

no matter what \(b\) is

OpenStudy (whovianchick):

Gotcha. So would be the answer for the first part of the question, f=g?

OpenStudy (whovianchick):

would 1*

satellite73 (satellite73):

so we know they are equal if \(x=1\) since you get 0 for both

satellite73 (satellite73):

yes

satellite73 (satellite73):

and evidently if \(x>1\) you have \(\log_4(x)>\log_7(x)\) since you have to raise 4 to a bigger power

OpenStudy (whovianchick):

ok! And if x<1, then log 7(x) is bitter than log 4 (x)?

satellite73 (satellite73):

bitter and bigger too i think

satellite73 (satellite73):

they will of course both be negative

OpenStudy (whovianchick):

Ok, thanks!